# Paying Attention to Deflections: Mining Pragmatic Nuances for Whataboutism Detection in Online Discourse

### v1

@@ Abstract
Khiem Phi * Noushin Salek Faramarzi * Chenlu Wang * Ritwik Banerjee * * Department of Computer Science Institute for AI-Driven Discovery and Innovation Stony Brook University, New York, USA {kphi, nsalekfarama, chenlwang, rbanerjee}@cs.stonybrook.edu
@@ 1 Introduction
Whataboutism is the practice of deflecting criticism or avoiding an unfavorable issue by raising a different, more favorable matter, or by making a counter-accusation. Since its first use in 1974 zimmer2017roots, it has emerged as a common variation of the classical fallacy known as tu quoque (lit. “you also”) – attacking the opponent’s behavior or action for being inconsistent with their argument, thereby discrediting them. Despite significant work devoted to misinformation and propaganda, the detection of whataboutism has largely relied on the ease of tracking “what about” phrases. This, however, is informed by popular notions of the phenomenon, leading to a naïve linguistic treatment and subsequently, a neglect of the unique challenges to its detection. Many “what about” phrases do not, in fact, signal propagandist use (Fig. 1).
Whataboutism drives information disorder333Following the Council of Europe report by wardle2017information, ‘information disorder” includes dis/misinformation (falsehoods with/out the intention to mislead), as well as malinformation (e.g., hate speech). by derailing proper argumentation, instead of the relatively conspicuous act of misleading by directly lying. As epistemologist fallis2015disinformation argues, information disorder comprises two equally odious functions: (a) creating false beliefs, and (b) preventing the creation of belief in truth. Empirical efforts target the former through models for fact-checking and fake news detection guo2022survey. Whataboutism, on the other hand, is a prevalent and potent contributor to the latter as well. Given its deleterious effects on discourse and social cohesion, it is thus critical for scalable propaganda detection models to be informed by a deeper understanding of whataboutism so that valid argumentation is not conflated with propaganda.
@@ 2 Pragmatic flexibility
“What about” whataboutism. We utilize the theory of structured meaning approach krifka2001structured for a deeper examination beyond early datasets that appear to have conflated the two to a large extent dasanmartino2020semeval. Studies of the “what about” construct demonstrate extremely high semantic and pragmatic complexity, but has received little empirical attention bledin2021topicality. It can be used with a broad range of syntactic constituents, yielding diverse semantics. Further, it has often been framed in terms of an antecedent question under discussion (QUD) roberts2012information, wherein a diversion is achieved by establishing the prejacent as a referential topic while simultaneously raising a new question about its properties ebert2014unified. Thus, it is not the bigram “what about”, but the nature of the newly raised question, which determines whether whataboutism is being introduced in a discourse. To wit, Fig. 1 depicts valid argumentation employing this phrase, while an instance of whataboutism does not.
Is it always propaganda? Whether “what about” is explicit or paraphrased, it can suggest a solution, challenge a statement, add specification to prior discussion, or laterally redirect the QUD beaver2017questions; bledin2020resistance. This extreme pragmatic flexibility makes it difficult to distinguish valid argumentation from propagandist use. The problem is further compounded by the fact that the pragmatic distinctions coexist with lexical and semantic similarity. Fig. 2 illustrates how the same text may serve as (i) harmful deflection via accusation or diversion, which are characteristics of propaganda, or (ii) accommodation or “centering” of attention.555The former characterize propaganda, while the latter is a valid act that, incidentally, has immense utility in NLP tasks (e.g., anaphora resolution grosz1983providing; dekker1994predicate). Thus, approaches solely relying on syntax or semantics may not differentiate these opposing discourse maneuvers.
@@ 3 The Datasets
dasanmartino2019fine curated a corpus for propaganda detection, containing 76 instances of whataboutism, which serves as a valuable foundation for the study of propaganda. However, its size and assumption (that every occurrence is a propagandist use), limit its suitability for our study. We introduce the TQ collections,666To highlight that modern use of whataboutism surpasses the classical tu quoque fallacy, our datasets are dubbed TQ, with platform-specific subscripts serving as mnemonic aids. encompassing the social discourse around each instance. We hope this dataset, featuring ten times more labeled instances than the previous corpus, will spur deeper analyses of fallacies and inspire further research on the role of pragmatics in propaganda and misinformation.
@@
YouTube comments have not received much attention in empirical studies. Our first collection comprises YouTube comments due to a balanced user distribution across demographics statista2022gender; statista2022nations2; statista2022nations1; statista2022users and a less regulated environment, fostering diverse opinions mejova2012political. Targeting socially divisive topics prone to whataboutism, we formulate search queries for six such topics on YouTube. From the top five most viewed videos per query, we collect English comments along with the title, transcript, the number of up-votes, and the publisher information. This corpus, TQ, comprises 1,642 labeled comments from 17 videos across 6 topics, sorted by up-votes.
@@
Similar to TQ, our Twitter dataset TQ focuses on socially relevant topics that tend to elicit strong emotional responses. We gather English tweets and their replies 8 such topics. The original tweet offers the pragmatic context for identifying whataboutism in responses. While six topics overlap with TQ, the other two contain fewer tweets, so that our corpus can be used to analyze model performance on topics with limited data. For each topic, we gather tweets with significant engagement, filtering out threads with less than 200 messages. Within each thread, we exclude messages that do not directly respond to the original tweet, lack opinions or discursive content (e.g., “wow”), are too socially inappropriate, consist solely of emojis/emotions, or contain images/videos. In TQ, each datum comprises a tweet-reply pair, and the collection consists of 1,202 messages.
@@
We divide both datasets into training, validation, and testing sections, ensuring an even distribution of comments across topics in each section. Within each topic, comments are split into 80% training, 5% validation, and 15% testing. These portions are then combined across topics to form the training, validation, and test sets for the entire dataset. The labeled data distribution along these sets, as well as the class-wise distributions, are described in Appendix LABEL:appendix:datasheet.
@@
We enlist three annotators with native fluency in English who independently assign binary labels to each comment.777shoukri2003measures suggests no additional benefits from more than three annotators for a dichotomous variable. To ensure context awareness, annotators are required to view the pertinent YouTube video or read the original tweet before labeling each instance. The final label is determined by majority vote, with inter-annotator agreements at (TQ) and (TQ).888There is no single threshold for a good value of . The widely used demarcations introduced by landis1977 regard these scores as substantial agreement.
@@
Following recent recommendations gebru2018datasheets; bender2018data, we include two comprehensive transparency artifacts: (a) the data statement (Appendix LABEL:appendix:datasheet) and (b) the annotation guide/codebook (included in our data repository).
@@
YT Video: One dead after ‘Unite the Right’ rally in Virginia (Fox News) |
(1) if such things happened in china and russia, what would cnn and nytimes say? |
(2) If the perp was an Islamist, imagine how the left would dismiss what he did as not being indicative of Islam (like they always do). |
[…]
(3) They did the same thing to indigenous folks for bounty hunters and military, the US military deserves this. |
[…]
(4) Fox News, Republicans and Trump all defended this monster. They are now hoping and praying that you forget that part. |
[…]
(5) Russia bombing buildings four miles from Kyiv’s center doesn’t sound like losing. How close have the Ukrainians gotten to Moscow? |
(6) USA baby not didn’t war crime afganistan Iraq |
[…]
(7) Coming from a govt that congratulated Bola Tinubu of Nigeria. Funny people |
(8) And what of Ukraine’s war crimes? Torture and murder of POWs, firing from civilian positions, using human shields or shooting civilians for receiving Russian ration packs. ALL war crimes need to be prosecuted, not just those by Russians. |
@@ 3.1 A thematic scrutiny
YouTube comment-threads lack coherent discourse among users, with each comment typically serving as a standalone response to the video or earlier comments. This differs from discourse structures in articles, interviews, or debates, where whataboutism is commonly observed putz2016whataboutism; dykstra2020rhetoric. While Twitter threads also show some deviation, our filtering steps ensure that TQ maintains a cleaner statement-response format.
@@
The users seldom employ whataboutism as a defensive tactic to deflect criticism. Instead, it is frequently wielded to preclude the relevant issue by introducing an accusation, diverging from the traditional view of whataboutism as a response with a counter-accusation. Users frequently employ whataboutism to shape perceptions, understanding, or evaluation of an issue, particularly through the lens of domestic political divides. Table 3 illustrates this trend: comments on a Fox News video critique liberal media outlets like CNN and the New York Times (1, 2), while criticisms of conservative entities appear under content from liberal organizations (4), and even centrist channels (3).999The liberal, conservative, or centrist leanings mentioned in this discussion are obtained from adfontes.
@@
Users commonly employ whataboutism to highlight perceived selection bias or hypocrisy in news coverage, reframing overlooked issues in a new context. Assessing whether such reframing constitutes propaganda is complex, given its prevalence even in journalism from reputable organizations.101010For example: (1) Dhanesha, N. Climate fixes are all aimed at property owners. What about renters? July 27, 2022. Vox. Accessed: 06.23.2023. (2) Hawkins, A. and Davidson, H. As the west tries to limit TikTok’s reach, what about China’s other apps? April 12, 2023. The Guardian. Accessed: 01.03.2024. Social media users similarly engage in this practice: in (5), a central claim is challenged with what the author views as overlooked evidence;111111Table 3:(5) implies there are the only possible outcomes, showcasing the false dichotomy fallacy. and (7) demonstrates the use of whataboutism in an appeal to hypocrisy but deflecting not as a counter-accusation, diverging from tu quoque.
@@
There are similarities between whataboutism in propaganda, whataboutism in general discourse, and the “what about” bigram, but our thematic analysis reveals distinct characteristics, preventing one from implying the other, despite topical redirection. Some comments, like Table 3:(4), overtly redirect without using whataboutism, while others, like (5), attempt to reframe the same topic. Furthermore, contemporary whataboutism differs from the classical fallacy as they deflect to entities not directly involved in the conversation. Rather than a two-person interaction, commenters may perceive the video or previous posts as representative of “tu”. For instance, in Table 3:(3, 6), they may view the videos as representing the U.S. government and comment on past military actions. However, instances like (2) introduce accusations unrelated to the discourse (“the left”). Twitter examples, like (8), align more closely with classical tu quoque, likely due to the direct tweet-response structure.
Finally, since subjectivity and opinion influence people significantly picard2000affective, we analyze our datasets for subjective expressions and emotive content. We find that explicit sentiment polarity or emotions are rare, while sarcasm or irony abound, as seen in Table 3:(1, 6, 7).
@@ 4.1 Baselines: (Habitual) Transfer learning
To establish competitive baselines, we look to the best performing models from the SemEval-2020 propaganda detection task dasanmartino2020semeval. We add a fully connected layer and fine-tune them with these configurations on a single NVIDIA Titan XP GPU: (i) batch size: 80 for BERT-base and 40 for RoBERTa, (ii) 10 epochs, (iii) maximum sequence length: 256, and (iv) : .
Further, along the lines adopted by vlad2019sentence and yu2021interpretable, we incorporate a baseline that integrates affective language with task-specific fine-tuning. Due to our corpus being rich in irony and sarcasm, while being sparse in explicit expressions of sentiment, we fine-tune a RoBERTa model pretrained on irony detection barbieri2020tweeteval with the same training configuration.
@@
Identifying whataboutism hinges on discerning a sentence’s relation to the antecedent question under discussion (QUD), given its diversionary nature. For improvements over baseline results, we explore models that promise a deeper semantic understanding within a topic’s discourse. Given inadequate performance – particularly on YouTube comments – of language models pretrained on token prediction, we adopt SBERT reimers2019sentence, a model that is trained on natural language inference (NLI) datasets bowman2015large; williams2018broad, and excels in NLI and semantic textual similarity (STS) benchmarks. We fine-tune its embeddings with a configuration identical to that described earlier for BERT.
We also explore two additional contrastive learning methods: (i) “Correct and Smooth” (C&S) huang2021combining, which constructs a graph where the embeddings serve as nodes and cosine similarity provides edge weights, and tunes the embeddings via label propagation zhou2003learning; and (ii) SimCSE gao2021simcse, which has achieved notable results in STS tasks with supervision from annotated pairs in NLI datasets (entailment and contradiction pairs serving as positive and hard negative samples, respectively).
Notwithstanding their proficiency in semantic similarity benchmarks, our experiments reveal (§ 5) these models to be inadequate in understanding the nuanced pragmatic variations prevalent within each topic’s social discourse in our dataset. A common thread in the methodology of these models is their use of cosine similarity to differentiate between instances from positive and negative samples. Cosine similarity may present challenges in high-dimensional vector spaces, as it becomes highly probable that any two vectors are nearly orthogonal (see Appendix B for a proof). Further, as experimental results reveal significant room for improvement over these models, we conjecture that current approaches relying on cosine measure may struggle to detect whataboutism. This challenge drives us to explore beyond semantic similarity to capture pragmatic differences in modern social media discourse, with empirical evidence in § 5.
(a) Whataboutism detection on YouTube comments | ||||||
Model |
W | NW | ||||
P | R | F1 | P | R | F1 | |
Transfer learning (baseline models): | ||||||
bert | 0.72 | 0.36 | 0.48 |
0.92 |
0.98 |
0.95 |
roberta |
0.23 | 0.71 | 0.34 |
0.94 |
0.64 |
0.76 |
roberta Irony |
0.30 | 0.30 | 0.30 |
0.93 |
0.99 |
0.96 |
Based on conventional measures of semantic similarity: | ||||||
sbert | 0.13 | 0.05 | 0.07 |
0.96 |
0.99 |
0.97 |
c&s | 0.45 | 0.45 | 0.45 |
0.91 |
0.91 |
0.91 |
simcse |
0.27 | 0.68 | 0.38 |
0.94 |
0.72 |
0.82 |
Mining negatives with attention (mina): | ||||||
sbert | 0.63 | 0.53 | 0.58 |
0.94 |
0.96 |
0.95 |
roberta Irony |
0.57 | 0.47 | 0.52 |
0.94 |
0.95 |
0.94 |
(b) Whataboutism detection on Twitter replies | |||||||
Model |
W | NW | Params | ||||
P | R | F1 | P | R | F1 |
@@
bert | 0.64 | 0.76 | 0.69 |
0.85 |
0.76 |
0.80 |
110 |
roberta |
0.64 | 0.80 | 0.71 |
0.87 |
0.75 |
0.81 |
124 |
roberta Irony |
0.54 | 0.89 | 0.67 |
0.91 |
0.57 |
0.70 |
124 |
@@
sbert | 0.73 | 0.78 | 0.75 |
0.87 |
0.84 |
0.85 |
31 |
roberta Irony |
0.91 | 0.57 | 0.70 |
0.54 |
0.89 |
0.67 |
133 |

### v2

@@ Abstract
Khiem Phi Noushin Salek Faramarzi Chenlu Wang Ritwik Banerjee Department of Computer Science Stony Brook University, New York, USA {kphi, nsalekfarama, chenlwang, rbanerjee}@cs.stonybrook.edu
@@ 1 Introduction
Whataboutism is the practice of deflecting criticism or avoiding an unfavorable issue by raising a different, more favorable matter, or by making a counter-accusation. Since its first use in 1974 Zimmer (2017), it has emerged as a common variation of the classical fallacy known as tu quoque (lit. “you also”) – attacking the opponent’s behavior or action for being inconsistent with their argument, thereby discrediting them. Despite significant work devoted to misinformation and propaganda, the detection of whataboutism has largely relied on the ease of tracking “what about” phrases. This, however, is informed by popular notions of the phenomenon, leading to a naïve linguistic treatment and subsequently, a neglect of the unique challenges to its detection. Many “what about” phrases do not, in fact, signal propagandist use (Fig. 1).
Whataboutism drives information disorder333Following the Council of Europe report by Wardle and Derakhshan (2017, pp. 10, 20), ‘information disorder” includes dis/misinformation (falsehoods with/out the intention to mislead), as well as malinformation (e.g., hate speech). by derailing proper argumentation, instead of the relatively conspicuous act of misleading by directly lying. As epistemologist Fallis (2015, p. 420) argues, information disorder comprises two equally odious functions: (a) creating false beliefs, and (b) preventing the creation of belief in truth. Empirical efforts target the former through models for fact-checking and fake news detection Guo et al. (2022). Whataboutism, on the other hand, is a prevalent and potent contributor to the latter as well. Given its deleterious effects on discourse and social cohesion, it is thus critical for scalable propaganda detection models to be informed by a deeper understanding of whataboutism so that valid argumentation is not conflated with propaganda.
@@ 2 Pragmatic flexibility
“What about” whataboutism:
We utilize the theory of structured meaning approach Krifka (2001) for a deeper examination beyond early datasets that appear to have conflated the two to a large extent Da San Martino et al. (2020). Studies of the “what about” construct demonstrate extremely high semantic and pragmatic complexity, but has received little empirical attention Bledin and Rawlins (2021). It can be used with a broad range of syntactic constituents, yielding diverse semantics. Further, it has often been framed in terms of an antecedent question under discussion (QUD) Roberts (2012), wherein a diversion is achieved by establishing the prejacent as a referential topic while simultaneously raising a new question about its properties Ebert et al. (2014). Thus, it is not the bigram “what about”, but the nature of the newly raised question, which determines whether whataboutism is being introduced in a discourse. To wit, Fig. 1 depicts valid argumentation employing this phrase, while an instance of whataboutism does not.
Whether “what about” is explicit or not, it can suggest a solution, challenge a statement, add specification to prior discussion, or laterally redirect the QUD Beaver et al. (2017); Bledin and Rawlins (2020). This extreme pragmatic flexibility makes it difficult to distinguish valid argumentation from propagandist use. The problem is further compounded by the fact that the pragmatic distinctions coexist with lexical and semantic similarity. Fig. 2 illustrates how the same text may serve as (i) harmful deflection via accusation or diversion, which are characteristics of propaganda, and (ii) accommodation or “centering” of attention.555The former characterize propaganda, while the latter is a valid act that, incidentally, has immense utility in NLP tasks like anaphora resolution Grosz et al. (1983); Dekker (1994). Thus, strategies relying solely on syntax or semantics may not differentiate these opposing discourse maneuvers.
@@ 3 The Datasets
Da San Martino et al. (2019b) curated a corpus for propaganda detection, containing 76 instances of whataboutism, which serves as a valuable foundation for the study of propaganda. However, its size and assumption (that every occurrence is a propagandist use) limit its suitability for our study. We introduce the TQ collections,666To highlight that modern use of whataboutism surpasses the classical tu quoque fallacy, our datasets are dubbed TQ, with platform-specific subscripts serving as mnemonic aids. encompassing the social discourse around each instance. We hope this dataset, featuring ten times more labeled instances than the previous corpus, will spur deeper analyses of fallacies and inspire further research on the role of pragmatics in propaganda and misinformation.
@@
YouTube comments have not received much attention in empirical studies. Our first collection comprises YouTube comments due to a balanced user distribution across demographics Statista Research Department (2022a, b, c, d) and a less regulated environment, fostering diverse opinions Mejova and Srinivasan (2012). Targeting socially divisive topics prone to whataboutism, we formulate search queries for six such topics on YouTube. From the top five most viewed videos per query, we collect English comments along with the title, transcript, the number of up-votes, and the publisher information. This corpus, TQ, comprises 1,642 labeled comments from 17 videos across 6 topics, sorted by up-votes.
@@
Similar to TQ, our Twitter dataset TQ focuses on socially relevant topics that tend to elicit strong emotional responses. We gather English tweets and their replies 8 such topics. The original tweet offers the pragmatic context for identifying whataboutism in responses. While six topics overlap with TQ, the other two contain fewer tweets, so that our corpus can be used to analyze model performance on topics with limited data. For each topic, we gather tweets with significant engagement, filtering out threads with less than 200 messages. Within each thread, we exclude messages that do not directly respond to the original tweet, lack opinions or discursive content (e.g., “wow”), are socially too inappropriate, consist solely of emojis/emotions, or contain images/videos. In TQ, each datum comprises a tweet-reply pair, and the collection consists of 1,202 messages.
@@
We use stratified partitioning to divide both datasets into training, validation, and testing sections, ensuring an even distribution of comments across topics in each. Within each topic, comments were split into 80% training, 5% validation, and 15% testing, maintaining the class distribution throughout. These stratified segments are then combined across topics to form the final training, validation, and test sets. The distribution of labeled data across these sets, as well as the class-wise distributions, are described in Appendix A.
@@
We enlist three annotators with native fluency in English who independently assign binary labels to each comment.777Shoukri (2003) suggests no additional benefits from more than three annotators for a dichotomous variable. To attain an understanding of the broader sociopolitical discourse pertaining to each topic, annotators dedicate considerable time to carefully review the entire YouTube video or read the complete Twitter conversation before labeling each instance. This meticulous process is expected to ensure diligent annotation and achieve a high level of data fidelity, despite limiting the size of the datasets. The final label is determined by majority vote, with inter-annotator agreements measured by Fleiss’ kappa Fleiss and Cohen (1973) at (TQ) and (TQ).888There is no single threshold for a good value of . The widely used interpretations introduced by Landis and Koch (1977) regard these scores as “substantial agreement”.
@@
Following recent recommendations Gebru et al. (2021); Bender and Friedman (2018), we include two comprehensive transparency artifacts: (a) the data statement (Appendix A) and (b) the annotation guide/codebook, included in our data repository.
@@
YT Video: One dead after ‘Unite the Right’ rally in Virginia (Fox News) | (1) if such things happened in china and russia, what would cnn and nytimes say? | ✓ | ✓ | |
(2) If the perp was an Islamist, imagine how the left would dismiss what he did as not being indicative of Islam (like they always do). | ✓ | ✓ | ||
YT Video: Outrage grows over Russian bounties (ABC News) | (3) They did the same thing to indigenous folks for bounty hunters and military, the US military deserves this. | ✓ | ||
YT Video: Russians Flee Into Exile Because Of Putin’s War With Ukraine: NYT (MSNBC) | (4) Fox News, Republicans and Trump all defended this monster. They are now hoping and praying that you forget that part. | ✓ | ||
YT Video: Russia is aware it is not winning the war in Ukraine: Brookings senior fellow (CNBC) | (5) Russia bombing buildings four miles from Kyiv’s center doesn’t sound like losing. How close have the Ukrainians gotten to Moscow? | ✓ | ✓ | |
(6) USA baby not didn’t war crime afganistan Iraq | ✓ | ✓ | ✓ | |
Tweet: We united the world to protect Ukraine, we will unite the world to restore justice. Russian invaders will be legally and fairly held to account for all war crimes. The terrorist state will be held to account for the crime of aggression. (ZelenskyyUa; Mar 3, 2023) | (7) Coming from a govt that congratulated Bola Tinubu of Nigeria. Funny people | ✓ | ✓ | |
(8) And what of Ukraine’s war crimes? Torture and murder of POWs, firing from civilian positions, using human shields or shooting civilians for receiving Russian ration packs. ALL war crimes need to be prosecuted, not just those by Russians. | ✓ | ✓ | ✓ |
@@ 3.1 A thematic scrutiny
Inspired by the wide adoption of thematic analysis in qualitative research Braun and Clarke (2006); Guest et al. (2011), we offer a qualitative discussion to uncover recurring themes in our data. It stands not as an extension, but in complement, to the quantitative study based on corpus annotations. While rare in computational research, thematic analysis has been closely linked to foundational work on word senses and content analysis Stone et al. (1966); Litkowski (1997); Ide and Véronis (1998).
Discourse coherence:
@@
Comment-threads on YouTube videos lack coherent discourse. A comment often serves as a standalone response to the video or earlier comments. This differs from discourse structures in articles, interviews, or debates, where whataboutism has been observed Putz (2016); Dykstra (2020). While Twitter threads also show some deviation, our filtering ensures that TQ offers a cleaner statement-response format.
Perceptions, framing, and Tu Quoque:
@@
We see that users seldom employ whataboutism as a defensive tactic to deflect criticism. Instead, it is frequently wielded to preclude the relevant issue by introducing an accusation, diverging from the traditional view of whataboutism as a response with a counter-accusation. Users frequently employ whataboutism to shape perceptions, understanding, or evaluation of an issue, particularly through the lens of domestic political divides. Table 1 illustrates this trend: comments on a Fox News video critique liberal media outlets like CNN and the New York Times (1, 2), while criticisms of conservative entities appear under content from liberal organizations (4), and even centrist channels (3).999The liberal, conservative, or centrist leanings mentioned in this discussion are obtained from Ad Fontes Media (2022).
@@
Users commonly employ whataboutism to highlight perceived selection bias or hypocrisy in news coverage, reframing overlooked issues in a new context. Assessing whether such reframing constitutes propaganda is complex, given its prevalence even in journalism from reputable organizations.101010For example: (1) Dhanesha, N. Climate fixes are all aimed at property owners. What about renters? July 27, 2022. Vox. Accessed: 06.23.2023. (2) Hawkins, A. and Davidson, H. As the west tries to limit TikTok’s reach, what about China’s other apps? April 12, 2023. The Guardian. Accessed: 01.03.2024. Social media users similarly engage in this practice: in (5), a central claim is challenged with what the author views as overlooked evidence;111111Table 1:(5) implies there are the only possible outcomes, showcasing the false dichotomy fallacy. and (7) demonstrates the use of whataboutism in an appeal to hypocrisy but the deflection is not a counter-accusation, diverging from the tu quoque fallacy.
@@
There are similarities between whataboutism in propaganda, whataboutism in general discourse, and the “what about” bigram. Our thematic analysis, however, reveals distinct characteristics, preventing one from implying the other, despite topical redirection. Some comments, like Table 1:(4), overtly redirect without using whataboutism, while others, like (5), attempt to reframe the same topic. Furthermore, contemporary whataboutism differs from the classical fallacy as they deflect to entities not directly involved in the conversation. Rather than a two-person interaction, commenters may perceive the video or previous posts as representative of “tu”. For instance, in Table 1:(3, 6), they may view the videos as representing the U.S. government and comment on past military actions. However, instances like (2) introduce accusations unrelated to the discourse (“the left”). Twitter examples, like (8), align more closely with classical tu quoque, likely due to the direct tweet-response structure.
Sarcasm and irony:
@@ 4.1 Baselines: (Habitual) Transfer learning
To establish competitive baselines, we look to the best performing models from the SemEval-2020 propaganda detection task Da San Martino et al. (2020).121212While Piskorski et al. (2023) present a related task, we opt not to use the top models of that task due to: (a) the multilingual corpus; (b) their use of fine-tuned Transformer models (or their ensembles), much like the models in the earlier task; and (c) the poorer English-language whataboutism detection results, compared to the SemEval-2020 task. We incorporate a fully connected layer and fine-tune them using the following configurations on a single NVIDIA Titan XP GPU: (i) batch size of 80 for BERT-base and 40 for RoBERTa, (ii) 10 epochs, (iii) maximum sequence length of 256, and (iv) learning rate () set to .
Further, along the lines adopted by Vlad et al. (2019) and Yu et al. (2021), we include another baseline that integrates affective language with task-specific fine-tuning. Due to our corpus being rich in irony and sarcasm but sparse in explicit expressions of sentiment, we fine-tune a RoBERTa model pretrained on irony detection Barbieri et al. (2020) with the same training configuration.
@@
Given the diversionary nature of whataboutism, identifying it hinges on discerning a sentence’s relation to the antecedent question under discussion (QUD). For improvements over baseline results, we thus explore models that promise a deeper semantic understanding within a topic’s discourse. Given the inadequate performance – particularly on YouTube comments – of language models pretrained on token prediction, we adopt SBERT Reimers and Gurevych (2019), which is trained on natural language inference (NLI) datasets Bowman et al. (2015); Williams et al. (2018), and excels in NLI and semantic textual similarity (STS) benchmarks. We fine-tune its embeddings with a configuration identical to that described earlier for BERT.
We also explore two additional contrastive learning methods: (i) “Correct and Smooth” (C&S) Huang et al. (2021), which constructs a graph where the embeddings serve as nodes and cosine similarity provides edge weights, and tunes the embeddings via label propagation Zhou et al. (2003); and (ii) SimCSE Gao et al. (2021), which has achieved notable results in STS tasks with supervision from annotated pairs in NLI datasets (with entailment and contradiction pairs serving as positive and negative samples, respectively).
(a) Whataboutism detection on YouTube comments | |||||||
Model | W | NW | |||||
P | R | F1 | P | R | F1 | ||
@@
bert | 0.72 | 0.36 | 0.48 | 0.92 | 0.98 | 0.95 | |
roberta | 0.23 | 0.71 | 0.34 | 0.94 | 0.64 | 0.76 | |
roberta Irony | 0.30 | 0.30 | 0.30 | 0.93 | 0.99 | 0.96 | |
@@
sbert | 0.63 | 0.53 | 0.58∗ | 0.94 | 0.96 | 0.95 | |
roberta Irony | 0.57 | 0.47 | 0.52 | 0.94 | 0.95 | 0.94 |
(b) Whataboutism detection on Twitter replies | ||||||||
Model | W | NW | Params | |||||
P | R | F1 | P | R | F1 | () | ||
Transfer learning (baseline models): | ||||||||
bert | 0.64 | 0.76 | 0.69 | 0.85 | 0.76 | 0.80 | 110 | |
roberta | 0.64 | 0.80 | 0.71 | 0.87 | 0.75 | 0.81 | 124 | |
roberta Irony | 0.54 | 0.89 | 0.67 | 0.91 | 0.57 | 0.70 | 124 | |
Based on conventional measures of semantic similarity: | ||||||||
sbert | 0.64 | 0.76 | 0.69 | 0.85 | 0.76 | 0.80 | 22.7 | |
c&s | 0.57 | 0.60 | 0.59 | 0.95 | 0.91 | 0.93 | 22.7 | |
simcse | 0.64 | 0.80 | 0.71 | 0.87 | 0.74 | 0.80 | 22.7 | |
Mining negatives with attention (mina): | ||||||||
sbert | 0.73 | 0.78 | 0.75∗ | 0.87 | 0.84 | 0.85 | 31 | |
roberta Irony | 0.91 | 0.57 | 0.70 | 0.54 | 0.89 | 0.67 | 133 |
Notwithstanding their proficiency in STS benchmarks, the results of our experiments (§ 5) reveal these models to be inadequate in understanding the pragmatic variations prevalent within each topic’s social discourse. A common thread in their methodology is the use of cosine similarity to differentiate between instances from positive and negative samples. Cosine similarity may present challenges in high-dimensional vector spaces, as it becomes highly probable that any two vectors are nearly orthogonal (Appendix B offers a formal proof). Further, as experiments reveal significant room for improvement over these models, we conjecture that approaches relying on cosine measure may struggle to detect whataboutism. This challenge drives us to explore beyond semantic similarity to capture pragmatic differences in modern social media discourse, with empirical results described in § 5.